OpenStudy (anonymous):

x2 - 3x = 4 (the 2 is smal)

OpenStudy (lilg132):

x^2 - 3x - 4 = 0 what two numbers added together give you -3 and multiply to give you -4?

OpenStudy (anonymous):

A?

OpenStudy (lilg132):

LOL

OpenStudy (anonymous):

A) 1 4 B) 1 -4 C) -1 4 D) -1 -4

OpenStudy (mimi_x3):

Just use wolfram alpha if you are lazy

OpenStudy (lilg132):

i'm not giving you anymore answers if you dont want to learn

OpenStudy (anonymous):

i wanna learn just not too good at this lol :)

OpenStudy (anonymous):

\[(x-4)(x+1)=0\]

OpenStudy (lilg132):

well stop guessing and try and understand

OpenStudy (lilg132):

what two numbers can i add together to give me -3

OpenStudy (anonymous):

1+2

OpenStudy (lilg132):

no that will give you +3 not -3

OpenStudy (lilg132):

-4 + ? BIG CLUE

OpenStudy (anonymous):

1-2?

OpenStudy (lilg132):

that will give you -1

OpenStudy (anonymous):

1+2-6?

OpenStudy (lilg132):

i said two numbers

OpenStudy (lilg132):

not three

OpenStudy (anonymous):

0-3?

OpenStudy (lilg132):

try taking a look at where i said BIG CLUE !

OpenStudy (lilg132):

good your getting there

OpenStudy (lilg132):

that would give you -3 but we need to find two numbers that give you -3 and the same two numbers need to multiply to give you -4

OpenStudy (anonymous):

plz i only have 10 minutes left

OpenStudy (lilg132):

its touching if you want to learn you will make more time

OpenStudy (anonymous):

plz i want to learn but just dont have time right now

OpenStudy (anonymous):

ok so you have x2 - 3x = 4 the first thing you want to do is equal this to zero, we can do that by subtracting 4 x^2-3x-4=0 Now, like what that person said, what can you multiply to get -4, and add to get -3? now let's guess and check. 1 * 4 = 4, but 1 + 4 = 5, ew that won't work hmm. 1 * -4 = -4, and 1 - 4 = -3 that works! now we use these two numbers, 1 and -4 and plug it back into the original equation (take the -3x away) x^2+1x-4x-4 basically what we did is split the -3 apart, because you see, 1x-4x equals -3x so it's basically the same thing, but the point of doing this is that you can factor by grouping. so lets group the first two terms, x^2 and 1x, what two numbers, or variables, go into both of these? well, x does. x Now what's x divided by x^2? that's just x x ( x now what's x divided by 1x? 1. x ( x +1 ) you know this is correct when you use the distributive propetry and get what you use to have now, let's factor the other two terms, -4x and -4 these look like they both go into -4 -4 now, -4x divided by 4? just x. -4 ( x -4 divided by -4? 1. -4 (x+1) Now, lets bring it all together. the numbers in the parentheses should be the same (you got x+1 in both, so it's right. (x+1) now let's bring over the numbers that aren't in parentheses x and -4 just put those in parentheses like (x-4) the factored version is (x+1)(x-4) now, we have to equal both of these to zero and solve like a regular algebra problem to get your two solutions x+1=0 subtract 1 from both sides x=-1 do the same thing to the other one x-4=0 add 4 to both sides x=4 those are your two solutions x=-1 and x=4